# Emancipatory Information Retrieval

### v4

@@ Abstract
Our world today is facing a confluence of several mutually reinforcing crises each of which intersects with concerns of social justice and emancipation. This paper is a provocation for the role of computer-mediated information access in our emancipatory struggles. We define emancipatory information retrieval as the study and development of information access methods that challenge various forms of human oppression, and situates its activities within broader collective emancipatory praxis. The term "emancipatory" here signifies the moral concerns of universal humanization of all peoples and the elimination of oppression to create the conditions under which we can collectively flourish. To develop an emancipatory research agenda for information retrieval (IR), in this paper we speculate about the practices that the community can adopt, enumerate some of the projects that the field should undertake, and discuss provocations to spark new ideas and directions for research. We challenge the field of IR research to embrace humanistic values and commit to universal emancipation and social justice. We also invite scholars from fields such as human-computer interaction, information sciences, media studies, design, social sciences, humanities, democratic theory, and critical theory, as well as legal and policy experts, civil rights and social justice activists, and artists to join us in realizing this transformation. In this process, we must both imagine post-oppressive worlds, and reimagine the role of IR in that world and in the journey that leads us there.
@@ 1 Introduction
Our world today is facing a confluence of several mutually reinforcing crises pushing us universally, but notably not uniformly, towards precarity. The recent COVID-19 pandemic did not only result in over seven million deaths [89] and disrupt everyday lives globally; it also laid bare and further reinscribed existing political, economic, and social hierarchies that are saliently colonial and racialized [103], gendered [29, 86], casteist [97], and ableist [85]. Likewise, the dangers of recent and impending climate catastrophes as well as the discourse around mitigative actions are inextricably intertwined with concerns of racial [120, 140, 105, 119], gender [54, 1], and disability [38] justice. Increasing global wealth and income inequality [31] both mirrors historical dimensions of marginalization and is contributing to erosion of democracies worldwide [109]. Globally, we are still reckoning with the ongoing legacies of settler-colonialism [143] and present-day neocolonialism [101], recontesting women’s rights to bodily autonomy [17], and witnessing growing assaults on transgender people’s right to exist [50, 68].
The range of crises we face in the twenty-first century requires us to reckon with multiple intersecting and co-constituting processes of hierarchical organization of our society and as a consequence with the systemic dehumanization, marginalization, and oppression of those designated to its bottom rung. Therefore, any attempt at addressing the crises of our times must begin with recommitment to the principles of social justice, universal humanization of all peoples, and emancipatory struggles against social, economic, and political oppression. This paper is a provocation for the role of computer-mediated information access in these struggles. Our aim is to challenge the field of information retrieval (IR) research to reimagine itself as what it can (or, ought to) be if grounded in humanistic values and committed to universal emancipation and social justice, and to encourage the community to explicitly articulate and make conscious choices about the values and incentives that shapes our field of computing. This is also an invitation to scholars from fields such as human-computer interaction (HCI), information sciences, media studies, design, social sciences, humanities, democratic theory, and critical theory, as well as legal and policy experts, civil rights and social justice activists, and artists to join us to collectively reimagine radically new sociotechnical futures and realize desired transformations.
@@
IR, the field of computing that is concerned with the design and study of information access systems, such as web search and recommender systems, has its roots in library science and linguistics [65]. Early IR research was primarily concerned with the design of effective indexing mechanisms that enable fast pattern-based lookup of text from large corpora. The modern-day successors of these systems, however, go far beyond serving the function of simple lookup. Modern search systems observe and model collective behaviors of how their users interact with retrieved information to infer intent behind user queries, and to predict users’ future needs and actions. Some of these systems may proactively surface new information and breaking news, recommend commercial products and services, and even suggest activities and actions for the user to undertake. Conversational systems, such ChatGPT, [^1] https://chat.openai.com/ Microsoft Copilot, [^2] https://copilot.microsoft.com/ and Google Gemini [^3] https://gemini.google.com/, may be entrusted to summarize retrieved information in response to users’ search queries, effectively shifting the burden of assessing the relevance and trustworthiness of retrieved information and their contextual interpretation from the user to the statistical models underlying these systems. These systems affect what information and perspectives receive exposure which shapes consumer behavior, political discourse, and culture [57, 56, 60], and directly influence material outcomes for people [116, 40].
@@
Today’s search and recommender systems are, therefore, much more than passive lookup tools. These systems have become embedded in our online social discourse, both shaping and being shaped by our values, cultural identities, sociopolitical beliefs, and shared understanding of our environment and ourselves. We could choose to see this transformation of information access systems from passive lookup tools to agents of influence as the information reaching out and permeating every aspect of our individual and social lives—akin to how Grudin [58] described the transition of computers from applications in narrow specialist domains to their present-day ubiquitous presence in our lives as “the computer reaches out”.
@@
Importantly, these systems and platforms may privilege certain perspectives and ways of knowing over others that entrench social hierarchies and intensify oppression. Or, alternatively they could serve to raise up marginalized voices and challenge the status quo. Therefore, they represent sites of conflict between exploitation and liberation. In both historical and ongoing conflicts, we can observe struggles for control over both traditional media (e.g., newspapers, radio, and television) and online media (i.e., social media and other online information access platforms). For the oppressed, they serve as medium for building shared understanding and solidarity necessary for social movements and collective action. For the oppressor, they represent sites to apply force to censor information and perspectives and manipulate public opinion. How should we then reflect on and reconcile the implications of IR research and development on our communities and our collective liberation? In this paper, we assert that the field of IR would benefit from critically reflecting on how our work relates to current sociopolitical orders and our possible sociotechnical futures, and argues for reimagining IR research grounded explicitly in the adoption of humanistic and emancipatory values and aspirations.
@@ 2 What is emancipatory IR?
We define emancipatory IR as the study and development of information access methods that challenge various forms of human oppression, and situates its activities within broader collective emancipatory praxis. The term “emancipatory” here signifies the moral concerns—central to this field of study—of universal humanization of all peoples and the elimination of oppression to create the conditions under which we can collectively flourish. It recognizes information and access to information both as profoundly political, and enacts research and development activities—such as, formalization, theorization, design, experimentation, publishing, open sourcing, deployment, platform governance, and community building—in service of universal struggles against all forms of structural oppression. It rejects the techno-deterministic premise that there is a single pre-determined path forward for information access technology development, and challenges the community to employ reflexivity to uncover the deeply embedded values, incentives, and sociotechnical imaginaries that shape the field of IR. It recognizes the agency of the field and its contributing members in shaping the world, while respecting that the challenges and necessary interventions are both fundamentally sociotechnical in nature. Therefore, the theories and practices of this field must be co-developed with cross-disciplinary collaboration including with scholars from HCI, design, information sciences, media studies, social sciences, humanities, democratic theory, and critical theory, as well as with legal and policy experts, civil rights and social justice activists, and artists, among others.
@@
To develop an emancipatory IR research agenda, we can take inspiration from other scholarly fields that have incorporated humanistic values and emancipatory aspirations. Young et al. [148] surveyed the body of emancipation research in information science, and identified four components of emancipation prevalent in the literature: Agency (i.e., freedom to act), dialogue (i.e., freedom to express), inclusion (i.e., freedom to belong), and rationality (i.e., freedom to think). Wright [144] defines emancipatory social science as an intellectual enterprise seeking to “generate scientific knowledge relevant to the collective project of challenging various forms of human oppression”. Wright outlines a framework that enumerates three specific tasks for the field: (i) Systematic diagnosis and critique of the world as it exists, (ii) envisioning viable alternatives, and (iii) elaborating a theory of social transformation. According to Wright, emancipatory social science is “a theory of a journey from the present to a possible future: the diagnosis and critique of society tells us why we want to leave the world in which we live; the theory of alternatives tells us where we want to go; and the theory of transformation tells us how to get from here to there – how to make viable alternatives, achievable”.
@@
In human–computer interaction (HCI), Bardzell and Bardzell [11, 10] discuss emancipation under the heading of humanistic HCI, which they define as research and practices in the field that deploy humanistic epistemologies and methodologies. They note that emancipatory HCI is “a fundamental goal of virtually all humanistic HCI”. Related, several strands of HCI research incorporate anti-oppressive and emancipatory [118], feminist [12, 14, 16, 15, 13], queer [83, 79], postcolonial and decolonial [69, 44, 122, 6, 70], anti-racist [4], anti-casteist [127, 128], anti-ableist [141, 121], post-capitalistic [49, 25], and anarchist [76, 84] epistemologies, theories and practices.
@@
Similarly, in the fields of machine learning (ML) and artificial intelligence (AI) there are several works that call for anti-oppressive and emancipatory [74, 113], anti-fascist [90], decolonial [5, 96], anti-casteist [73, 112], and abolitionist [19, 9, 46, 142] perspectives and design interventions. In data science, D’ignazio and Klein [43] and Guyan [59] discuss how feminist and queer epistemologies can inform the field, and Monroe-White [98] adopted Wright’s framework [144] to propose an emancipatory agenda for data science to mitigate the harms to marginalized populations.
@@ 3 Practices, projects, and provocations
To develop an emancipatory research agenda for IR, we speculate about the practices that the community can adopt, enumerate some of the projects that the field should undertake, and discuss provocations to spark new ideas and directions for research. However, the list of practices, projects, and provocations discussed here are neither immutable nor complete. It is ultimately the responsibility of the community engaged in emancipatory research to regularly reflect on, produce evidence-based critique, challenge, and reshape its agenda to affect tangible progress in our collective struggles against dehumanization, oppression, and marginalization. It may be worth reemphasizing at this stage that due to the saliently sociotechnical nature of the challenges in this area, the practices, projects, and provocations too will rarely reside strictly in the domain of computing and technology and must be informed by perspectives from outside of the field of computer science and the lived experiences of marginalized peoples. We illustrate the practices, projects, and provocations described in this section in Figure 1.
@@ 3.1 Practices of emancipatory IR
We adapt Wright’s emancipatory social science framework [144] to develop a framework for the practices of emancipatory IR. Like Wright’s, our framework consists of diagnosing and critiquing, imagining viable alternative futures, and elaborating our theories of change.
@@
  • Lack of representation and diversity within the IR community leading to propping up colonial cisheteropatriarchal capitalist values and erasure of feminist, queer, decolonial, anti-racist, anti-casteist, anti-ableist, and abolitionist perspectives
@@
The second practice of emancipatory IR is to imagine and develop desirable, viable, and achievable alternatives to our IR technologies and ways of organizing our communities with the aim of addressing the harms and injustices identified in the diagnosis and critique stage. Mitra [93] argues that the IR community should explicitly articulate the sociotechnical imaginaries [71, 72] that influence the design and development of IR technologies and platforms. He enquires “what would IR systems look like if designed for futures informed by feminist, queer, decolonial, anti-racist, anti-casteist, anti-ableist, and abolitionist thoughts, and if the focus of IR research was not to prop up colonial cisheteropatriarchal capitalist structures but to dismantle them?” This is a call for us to imagine post-oppressive worlds, and the role of IR both in that world and in the journey that leads us there.
We intentionally do not speculate about any specific futures in this paper. To satisfy that ask IR needs spaces and processes for design futuring [52]. The goal of futuring here is not just to envision a plurality of alternative utopian and protopian futures, but also to build shared understanding, aspirations, and commitment towards emancipatory outcomes through that participatory process. After all, the act of exercising one’s imagination itself can be liberating, or as Benjamin [20] calls it “an invitation to rid our mental and social structures from the tyranny of dominant imaginaries”.
@@
Our final practice concerns with the development of theories of change [136, 24, 125, 138] that makes explicit how we aim to realize the alternative futures in the face of current realities and how our own work contributes towards those goals. This should be grounded in theories of power and co-developed with cross-disciplinary scholars, legal and policy experts, activists, artists, and others to realize real structural changes. This involves both developing new research agendas as well as experimenting with and realizing new arrangements within the IR community and critically reflecting on our relationships with other institutions (e.g., industry and government). To provide concrete examples, we next discuss potential projects that we believe emancipatory IR should invest in.

### v6

@@ Abstract
Our world today is facing a confluence of several mutually reinforcing crises each of which intersects with concerns of social justice and emancipation. This paper is a provocation for the role of computer-mediated information access in our emancipatory struggles. We define emancipatory information retrieval as the study and development of information access methods that challenge various forms of human oppression, and situates its activities within broader collective emancipatory praxis. The term "emancipatory" here signifies the moral concerns of universal humanization of all peoples and the elimination of oppression to create the conditions under which we can collectively flourish. To develop an emancipatory research agenda for information retrieval (IR), in this paper we speculate about the practices that the community can adopt, enumerate some of the projects that the field should undertake, and discuss provocations to spark new ideas and directions for research. We challenge the field of IR research to embrace humanistic values and commit to universal emancipation and social justice. We also invite scholars from fields such as human-computer interaction, information sciences, media studies, design, science and technology studies, social and political sciences, philosophy, law, environmental sciences, public health, educational sciences, as well as legal and policy experts, civil rights advocates, social justice activists and movement organizers, and artists to join us in realizing this transformation. In this process, we must both imagine post-oppressive worlds, and reimagine the role of IR in that world and in the journey that leads us there.
Keywords: IR and society, Emancipatory praxis, Technology and power
@@ 1 Introduction
Our world today is facing a confluence of several mutually reinforcing crises pushing us universally, but notably not uniformly, towards precarity. The recent COVID-19 pandemic did not only result in over seven million deaths (mathieu2024coronavirus) and disrupt everyday lives globally; it also laid bare and further reinscribed existing political, economic, and social hierarchies that are saliently colonial and racialized (papamichail2023reinscribing), gendered (carli2020women; madgavkar2020covid), casteist (mondal2024caste), and ableist (lund2022ever). Likewise, the dangers of recent and impending climate catastrophes as well as the discourse around mitigative actions are inextricably intertwined with concerns of racial (sultana2024confronting; williams2022climate; patnaik2020racial; stefani2021indigenous), gender (gender2016gender; women2022explainer), and disability (cram2022cripping) justice. Increasing global wealth and income inequality (chancel2022world) both mirrors historical dimensions of marginalization and is contributing to erosion of democracies worldwide (rau2025income). Globally, we are still reckoning with the ongoing legacies of settler-colonialism (wolfe2006settler) and present-day neocolonialism (nkrumah1965neo), recontesting women’s rights to bodily autonomy (beck2024reproductive), and witnessing growing assaults on transgender people’s right to exist (human2023epidemic; horbury2020empire).
The range of crises we face in the twenty-first century requires us to reckon with multiple intersecting and co-constituting processes of hierarchical organization of our society and as a consequence with the systemic dehumanization, marginalization, and oppression of those designated to its bottom rung. Therefore, any attempt at addressing the crises of our times must begin with recommitment to the principles of social justice and universal humanization of all peoples and support for emancipatory struggles against social, economic, political, cultural, religious, sexual, and bodily oppression. This paper is a provocation for the role of computer-mediated information access in these struggles. Our aim is to challenge the field of information retrieval (IR) research to reimagine itself as what it can (or, ought to) be if grounded in humanistic values and committed to universal emancipation and social justice, and to encourage the community to explicitly articulate and make conscious choices about the values and incentives that shapes our field of computing. This is also an invitation to scholars from fields such as human-computer interaction (HCI), information sciences, media studies, design, science and technology studies (STS), social and political sciences, philosophy, law, environmental sciences, public health, educational sciences, as well as legal and policy experts, civil rights advocates, social justice activists and movement organizers, and artists to join us to collectively reimagine radically new sociotechnical futures and realize desired transformations.
@@
IR, the field of computing that is concerned with the design and study of information access systems, such as web search and recommender systems, has its roots in library science and linguistics (harman2019information). Early IR research was primarily concerned with the design of effective indexing mechanisms that enable fast pattern-based lookup of text from large corpora. The modern-day successors of these systems, however, go far beyond serving the function of simple lookup. Modern search systems observe and model collective behaviors of how their users interact with retrieved information to infer intent behind user queries, and to predict users’ future needs and actions. Some of these systems may proactively surface new information and breaking news, recommend commercial products and services, and even suggest activities and actions for the user to undertake. Conversational systems, such ChatGPT, [^1] https://chat.openai.com/ Microsoft Copilot, [^2] https://copilot.microsoft.com/ and Google Gemini [^3] https://gemini.google.com/, may be entrusted to summarize retrieved information in response to users’ search queries, effectively shifting the burden of assessing the relevance and trustworthiness of retrieved information and their contextual interpretation from the user to the statistical models underlying these systems. These systems affect what information and perspectives receive exposure which shapes consumer behavior, political discourse, and culture (grimmelmann2008google; gillespie2019algorithmically; hallinan2016recommended), and directly influence material outcomes for people (singh2018fairness; datta2014automated).
@@
Today’s search and recommender systems are, therefore, much more than passive lookup tools. These systems have become embedded in our online social discourse, both shaping and being shaped by our values, cultural identities, sociopolitical beliefs, and shared understanding of our environment and ourselves. We could choose to see this transformation of information access systems from passive lookup tools to agents of influence as the information reaching out and permeating every aspect of our individual and social lives—akin to how grudin1990computer described the transition of computers from applications in narrow specialist domains to their present-day ubiquitous presence in our lives as “the computer reaches out”.
@@
Importantly, these systems and platforms may privilege certain perspectives and ways of knowing over others as forms of epistemic injustice (fricker2007epistemic) that entrench social hierarchies and intensify oppression. Or, alternatively they could serve to raise up marginalized voices and challenge the status quo. Access to information is critical to collective sense-making of our place and relationships in this world. Therefore, it is unsurprising that throughout history authoritarian forces have tried to control what information is disseminated and how. In both historical and ongoing conflicts, we can observe these struggles for control over both traditional media (e.g., newspapers, radio, and television) and online media (i.e., social media and other online information access platforms). For the oppressed, they serve as medium for building shared understanding and solidarity necessary for social movements and collective action. For the oppressor, they represent sites to apply force to censor information and perspectives and manipulate public opinion. Information access platforms, therefore, represent sites of conflict between exploitation and liberation. How should we then reflect on and reconcile the implications of IR research and development on our communities and our collective liberation? In this paper, we assert that the field of IR would benefit from critically reflecting on how our work relates to current sociopolitical orders and our possible sociotechnical futures, and argues for reimagining IR research grounded explicitly in the adoption of humanistic and emancipatory values and aspirations.
@@ 2 What is emancipatory IR?
We define emancipatory IR as the study and development of information access methods that challenge all forms of human oppression, and situates its activities within broader collective emancipatory praxis. The term “emancipatory” here signifies the moral concerns—central to this field of study—of universal humanization of all peoples and the elimination of oppression to create the conditions under which we can collectively flourish. Historically, the term “emancipation” has taken on various meanings in context of diverging intellectual traditions (susen2015emancipation). It described the abolitionist struggles against slavery, but later came to signify a broader vision of dismantling all forms of structural oppression. Examples of structural oppression in this context include colonialism, racism, patriarchy, casteism, transphobia, religious persecution, and ableism. Emancipatory thinking in humanities have manifested in many forms, including Marxism, Critical Theory, feminism, and decolonialism. Emancipatory IR aims to be informed by and incorporate these different theories and epistemic frameworks in the conception and design of information access methods and systems.
This framing recognizes information and access to information both as profoundly political, and enacts research and development activities—such as, formalization, theorization, design, experimentation, publishing, open sourcing, deployment, platform governance, and community building—in service of universal struggles against all forms of structural oppression. It rejects the techno-deterministic premise that there is a single pre-determined path forward for information access technology development, and challenges the community to employ reflexivity to uncover the deeply embedded values, incentives, and sociotechnical imaginaries that shape the field of IR. It recognizes the agency of the field and its contributing members in shaping the world, while respecting that the challenges and necessary interventions are both fundamentally sociotechnical in nature. Therefore, the theories and practices of this field must be co-developed with cross-disciplinary collaboration including with scholars from HCI, design, information sciences, media studies, STS, social and political sciences, philosophy, law, environmental sciences, public health, educational sciences, as well as with legal and policy experts, civil rights advocates, social justice activists and movement organizers, and artists, among others.
Emancipatory IR challenges us to employ technological research and development in service of dismantling hierarchies of oppression and challenge power. It discourages non-performative academic gaze, and urges this research to be situated in movement building, and calls for recognizing the role of movement building practices within this research. It encourages us to prioritize praxis—i.e., research activities and reflections directed at structural change—over proxy metrics of success—e.g., state-of-the-art performances and leaderboard rankings that do not translate to scientific or social progress). And above all, it challenges the field to move beyond the restrictive view of IR research that emphasizes algorithmic advances as measurable on shared benchmarks and leaderboards to a more expansive view of IR research situated in a project of affecting social good and universal emancipation.
@@
To develop an emancipatory IR research agenda, we can take inspiration from other scholarly fields that have incorporated humanistic values and emancipatory aspirations. young2021emancipation surveyed the body of emancipation research in information science, and identified four components of emancipation prevalent in the literature: Agency (i.e., freedom to act), dialogue (i.e., freedom to express), inclusion (i.e., freedom to belong), and rationality (i.e., freedom to think). wright2020envisioning defines emancipatory social science as an intellectual enterprise seeking to “generate scientific knowledge relevant to the collective project of challenging various forms of human oppression”. Wright outlines a framework that enumerates three specific tasks for the field: (i) Systematic diagnosis and critique of the world as it exists, (ii) envisioning viable alternatives, and (iii) elaborating a theory of social transformation. According to Wright, emancipatory social science is “a theory of a journey from the present to a possible future: the diagnosis and critique of society tells us why we want to leave the world in which we live; the theory of alternatives tells us where we want to go; and the theory of transformation tells us how to get from here to there – how to make viable alternatives, achievable”.
@@
In HCI, bardzell2016humanistic; bardzell2015humanistic discuss emancipation under the heading of humanistic HCI, which they define as research and practices in the field that deploy humanistic epistemologies and methodologies. They note that emancipatory HCI is “a fundamental goal of virtually all humanistic HCI”. Related, several strands of HCI research incorporate epistemologies, theories and practices that are anti-oppressive and emancipatory (smyth2014anti), feminist (bardzell2010feminist; bardzell2011towards; bardzell2011feminism; bardzell2016feminist; bardzell2018utopias), queer (light2011hci; klipphahn2024introduction), postcolonial and decolonial (irani2010postcolonial; dourish2012ubicomp; sun2013critical; akama2016speculative; irani2016stories), anti-racist (abebe2022anti), anti-casteist (vaghela2022interrupting; vaghela2022caste), anti-ableist (williams2021articulations; sum2024challenging), post-capitalistic (feltwell2018grand; browne2022future), and anarchist (keyes2019human; linehan2014never).
@@
Similarly, in the fields of machine learning (ML) and artificial intelligence (AI) there are several works that call for perspectives and design interventions that are anti-oppressive and emancipatory (kane2021avoiding; saxena2023artificial), anti-fascist (mcquillan2022anti), decolonial (adams2021can; mohamed2020decolonial), anti-casteist (kalyanakrishnan2018opportunities; sambasivan2021re), and abolitionist (benjamin2019race; barabas2020beyond; earl2021towards; williams2023no). In data science, d2020data and guyan2022queer discuss how feminist and queer epistemologies can inform the field, and monroe2021emancipatory adopted Wright’s framework (wright2020envisioning) to propose an emancipatory agenda for data science to mitigate the harms to marginalized populations.
Towards emancipatory research in IR.
belkin1976some acknowledged information science’s social responsibility nearly half a century ago. But it was not until more recently that this perspective gained serious traction within the IR community, starting with the SWIRL (culpepper2018research) and the FACTS-IR (olteanu2021facts) workshops. Subsequently, there has been a wide range of IR research on fairness (ekstrand2021fairness), explainability (anand2022explainable), and addressing misinformation (zhou2020survey) among other societally-motivated topics. However, the field has been largely reluctant to acknowledge the saliently political nature of this work leaving the underlying colonial, cisheteropatriarchal, and capitalist values that has (and continue to) critically shape the field of IR unchallenged. Even IR conference tracks dedicated to research that tries to affect social good have often side-stepped the deeply sociopolitical question of defining what constitutes social good (mitra2025ir).
This has led to several recent calls to explicate and critically examine the norms and values (vrijenhoek2023report; trippas2025report) shaping the field as well as the sociotechnical futures that the IR community wants to realize through their research (mitra2025search; azzopardi2024report); while also reasserting the interdisciplinary roots of IR (zangerle2025beyond). This year, the ECIR IR-for-Good track’s call for papers [^4] https://ecir2026.eu/calls/call-for-ir-for-good-papers explicitly define IR-for-Good as “IR research and practices that contribute towards realizing more equitable, emancipatory, and sustainable futures” (mitra2025ir), and updated the list topics relevant to the track to shift the emphasis from desired attributes and interventions in IR systems (e.g., fairness and accessibility)—that were the norm in previous years [^5] https://ecir2025.eu/call-for-ir-for-good-papers/—to desired justice-oriented real-world outcomes (e.g., racial, disability, and sexuality justice) that IR research should try to realize. They also require papers to include explicit discussion on their theory of change.
Several of these developments—i.e.trippas2025report; azzopardi2024report; mitra2025ir—were prompted by conversations based on an earlier preprint of this current manuscript. In this work, we argue that moving forward the IR community should adopt emancipation as an explicit objective for the field.
@@ 3 Practices, projects, and provocations
To develop an emancipatory research agenda for IR, we speculate about the practices that the community can adopt, enumerate some of the projects that the field should undertake, and discuss provocations to spark new ideas and directions for research. However, the list of practices, projects, and provocations discussed here are neither immutable nor complete. It is ultimately the responsibility of the community engaged in emancipatory research to regularly reflect on, produce evidence-based critique, challenge, and reshape its agenda to affect tangible progress in our collective struggles against dehumanization, oppression, and marginalization. It may be worth reemphasizing at this stage that due to the saliently sociotechnical nature of the challenges in this area, the practices, projects, and provocations too will rarely be purely technological or reside strictly in the domain of computing, and must be informed by perspectives from outside of the field of computer science and the lived experiences of marginalized peoples. We illustrate the practices, projects, and provocations described in this section in Figure 1.
@@ 3.1 Practices of emancipatory IR
We adapt Wright’s emancipatory social science framework (wright2020envisioning) to develop a framework for the practices of emancipatory IR. Like Wright’s, our framework consists of diagnosing and critiquing, imagining viable alternative futures, and elaborating our theories of change.
@@
  • Lack of representation and diversity within the IR community leading to propping up colonial, cisheteropatriarchal, and capitalist values and erasure of feminist, queer, decolonial, anti-racist, anti-casteist, anti-ableist, and abolitionist perspectives
Diagnose and critique must extend to the policy and regulatory landscape, such as analyzing the European Union’s General Data Protection Regulation (GDPR)  (regulation2016regulation) and Digital Services Act (european2022regulation) in the context of our emancipatory goals which we identify as critical future work.
@@
The second practice of emancipatory IR is to imagine and develop desirable, viable, and achievable alternatives to our IR technologies and ways of organizing our communities with the aim of addressing the harms and injustices identified in the diagnosis and critique stage. Previous work (mitra2025search) have argued that the IR community should explicitly articulate the sociotechnical imaginaries (jasanoff2009containing; jasanoff2015dreamscapes) that influence the design and development of IR technologies and platforms; and enquires “what would IR systems look like if designed for futures informed by feminist, queer, decolonial, anti-racist, anti-casteist, anti-ableist, and abolitionist thoughts, and if the focus of IR research was not to prop up colonial cisheteropatriarchal capitalist structures but to dismantle them?” This is a call for us to imagine post-oppressive worlds, and the role of IR both in that world and in the journey that leads us there.
We intentionally do not speculate about any specific futures in this paper. To satisfy that ask, IR needs spaces and processes for design futuring (fry2009design). The goal of futuring here is not just to envision a plurality of alternative utopian and protopian futures, but also to build shared understanding, aspirations, and commitment towards emancipatory outcomes through that participatory process. After all, the act of exercising one’s imagination itself can be liberating, or as benjamin2024imagination calls it “an invitation to rid our mental and social structures from the tyranny of dominant imaginaries”.
@@
Our final practice concerns with the development of theories of change (weiss1995nothing; brest2010power; taplin2012theory; wiki:TheoryOfChange) that makes explicit how we aim to realize the alternative futures in the face of current realities and how our own work contributes towards those goals. This should be grounded in theories of power and co-developed with cross-disciplinary scholars, legal and policy experts, activists, artists, and others to realize real structural changes. This involves both developing new research agendas as well as experimenting with and realizing new arrangements within the IR community and critically reflecting on our relationships with other institutions (e.g., industry and government). To provide concrete examples, we next discuss potential projects that we believe emancipatory IR should invest in.